OpenStudy (anonymous):

A water hose 2cm in diameter is used to fil 20ltr bucket. If it takes 1 minute to fill the bucket, then the speed v at which the water leaves the hose is,(Ans: 106cm/s) Plz help me solving this problem Thank you.

OpenStudy (anonymous):

Inner area of hose = \[\pi r^2 = \pi \times 1^2 = 3.142cm^2\] 20 litres = \[20000 cm^3\] The water must be leaving the hose at \[20000/60 = 333.33 cm^3/second\] Therefore the speed in which the water must be leaving the hose is \[333.33 / 3.142 = 106 cm/s\] Hope this helps.

OpenStudy (anonymous):

Thank you